OpenStudy (anonymous):

use the wronskian to show that the given function are linearly independent on the given interval a. {1,x,x^2} on (negative infinity, positive infinity) b. {sinx,cosx,tanx} on ( negative pi over two, pi over two)

OpenStudy (turingtest):

the wronskian is\[W=\left|\begin{matrix}u&v&w\\u'&v'&w'\\u''&v''&w''\end{matrix}\right|\]which if it is nonzero, proves the linear independence of the solutions u, v, and w

OpenStudy (anonymous):

i tried the first one and i got a zero. so how can i prove its linearly independent?

OpenStudy (turingtest):

it isn't zero u=1 v=x w=x^2 so\[W=\left|\begin{matrix}u&v&w\\u'&v'&w'\\u''&v''&w''\end{matrix}\right|=\left|\begin{matrix}1&x&x^2\\0&1&2x\\0&0&2\end{matrix}\right|\neq0\]

OpenStudy (anonymous):

you are right sorry i missed up lol thank you so much. how about for the second one i get a very complex trig function. do i have to plug in the values of the interval ?

OpenStudy (turingtest):

you should get something that can be simplified, though it may be a bit tricky, to something that is never zero in that interval the way you wrote it the interval is open, so would not include those endpoints I haven't tried it...

OpenStudy (turingtest):

\[W=\left|\begin{matrix}u&v&w\\u'&v'&w'\\u''&v''&w''\end{matrix}\right|=\left|\begin{matrix}\sin x&\cos x&\tan x\\\cos x&-\sin x&\sec^2 x\\-\sin x&-\cos x&2\sec^2x\tan x\end{matrix}\right|\]\[=-\sin^2x(2\sec^2x\tan x)-\sec x\sin x-\cos^2x\tan x-\sin^2x\tan x...\]\[...+\cos x\sec^2 x\sin x-2\cos^2 x\sec^2x\tan x\]let's see if we can simplify....

OpenStudy (turingtest):

multiple typos :S\[=-\sin^2x(2\sec^2x\tan x)-\sec^2x\sin x\cos x-\cos^2x\tan x-\sin^2x\tan x...\]\[...+\cos x\sec^2 x\sin x-2\cos^2 x\sec^2x\tan x\]I hope it's right now...

OpenStudy (turingtest):

\[=-2\tan^3x-\tan x-\cos^2x\tan x-\sin^2x\tan x+\tan x-2\tan^3x\]hope I didn't mess up yet...

OpenStudy (turingtest):

\[W=-4\tan^3x-\tan x\]hmm... that has a point where it is zero I think I made some mistake somewhere :/

OpenStudy (anonymous):

first determinant has a sign error i think...+sec^2(x)sin(x)cos(x)

OpenStudy (anonymous):

\[i got -2\tan ^{3}x +\tan x -\cos x \sin x \sec ^{2}x-\tan x-\sin x \cos x+ \tan x \sin ^{2}x\]

OpenStudy (turingtest):

@mathmuse mine has that sign

OpenStudy (turingtest):

oh well, good luck finding my mistake I'm out !

OpenStudy (anonymous):

haha thanks anyways

OpenStudy (anonymous):

hmmm \[-2\tan^3x-2tanx-sinxcosx-\sin^2x\]

OpenStudy (anonymous):

either it tends to zero at x = 0.